OpenStudy (anonymous):

(x – 4 )(3x + 2) = 0 Solve for x, show work.

myininaya (myininaya):

since if we have a*b=0, we have a=0 or b=0

OpenStudy (anonymous):

first: FOIL it (first outer inner last) x times 3x x times 2 -4 times 3x -4 times 2

myininaya (myininaya):

set both factors equal to zero

myininaya (myininaya):

x-4=0 or 3x+2=0

myininaya (myininaya):

x=4 or 3x=-2

myininaya (myininaya):

x=4 or x=-2/3

OpenStudy (anonymous):

myininaya is correct.. do not foil this....

OpenStudy (anonymous):

now you should have 3x^2 + 2x - 12x - 8 = 0 from here combine like terms now you should have 3x^2 - 10x - 8 = 0